OpenStudy (yeval76):

***FAN AND MEDAL*** Erika is working on solving the exponential equation 50x = 17; however, she is not quite sure where to start. Using complete sentences, describe to Erika how to solve this equation.

OpenStudy (hasankoush):

If you're solving the equation for X then try to solve for it by separating X from 50, how would Erika do that?

OpenStudy (yeval76):

By turning it into a logarithmic function... i think @HasanKoush

jhonyy9 (jhonyy9):

,,exponential equation" - where is the term with exponent i dont see it anywhere ?

OpenStudy (yeval76):

Im sorry, it is \[50^x\]

jhonyy9 (jhonyy9):

ohhhh yes - this is different sure - yes ?

OpenStudy (yeval76):

yes

jhonyy9 (jhonyy9):

so and you know how you can start it to solve ?

OpenStudy (yeval76):

No thats why I am asking you... We are supposed to solve \[50^x = 17\]

OpenStudy (yeval76):

But I am not sure how to

jhonyy9 (jhonyy9):

opinion please

jhonyy9 (jhonyy9):

do you have anything ?

jhonyy9 (jhonyy9):

how you have learned it in your schoole ?

OpenStudy (yeval76):

i think that you have to first change each side into a log function but then after that I'm lost as to how to solve it

OpenStudy (mathmale):

You have 50^x = 17 and want to solve for x. What to do? Take the log or ln of both sides to eliminate the exponential function 50^x. What is log 50^x? What is log 17?

OpenStudy (yeval76):

In what I'm thinking, it would end up like this: \[xlog50 = \log 17\]

jhonyy9 (jhonyy9):

without knowladge about so you not can solve it ok try to make log in both sides - what you will get so ?

OpenStudy (yeval76):

\[xlog50 = \log 17\]

jhonyy9 (jhonyy9):

yes how you have wrote above is right - now just divide both sides with log 50 and will get the value of x - yes ?

OpenStudy (yeval76):

x = log 50 / log 17

jhonyy9 (jhonyy9):

no this is not rieght how you have made it - check it newly please

OpenStudy (yeval76):

Oh ok that was my question! Thank you so much, i was very confused... Do you think you could help me with another one? @mathmale

OpenStudy (mathmale):

x log 50 = log 17 has log17 on the right side. divide both sides of this equation by log 50 to obtain x = (log 17) / (log 50). End.

OpenStudy (yeval76):

Do you think you could help me with another one? @mathmale

OpenStudy (mathmale):

Please go ahead with your next question, as a new post.

OpenStudy (yeval76):

Ok